# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ 2.2 PEP for optimization with inexact oracle
and the Gram matrix $\mathbf{G}\triangleq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{S}^{2K+2}$ which is a symmetric positive semidefinite (PSD) matrix. Also, we define standard basis vectors $\vec{\mathbf{f}}_{k}\in\mathbb{R}^{K+1}$ and $\vec{\mathbf{x}}_{k},\vec{\mathbf{g}}_{k},\vec{\mathbf{e}}_{k}\in\mathbb{R}^{2K+2}$ to select corresponding columns in $\mathbf{F}$ or $\mathbf{X}$ to recover the needed element, which satisfy
@@
Note. We note that the subscript $k$ on the standard basis vectors does not mean that the $k$-th element of that vector is equal to 1. For instance, based on the definitions of $\mathbf{F}$ and $\mathbf{X}$ in (7), $\vec{\mathbf{f}}_{k}$ is a standard basis vector with $k+1$-th element equal to 1, $\vec{\mathbf{e}}_{k}$ is a standard basis vector with $k+2$-th element equal to 1, $\vec{\mathbf{g}}_{k}$ is a standard basis vector with $K+k+2$-th element equal to 1, and $\vec{\mathbf{x}}_{0}$ is a standard basis vector with the first element equal to 1. With a slight abuse of the notation, $\vec{\mathbf{x}}_{k}$ (which is not a standard basis vector) is the vector used to represent $\mathbf{x}_{k}$ by a linear combination of the columns of $\mathbf{X}$ and it is a function of $\theta_{k,i}$.
@@
The optimal solution of (SDP-PEP) will provide an exact worst-case performance of a given algorithm for convex problems. It is also possible to find the optimal algorithm with parameter ${\boldsymbol{\theta}}$ if we consider the minimax problem as
$\min_{\theta_{k,i}}\max_{\mathbf{G},\mathbf{F}\in\mathcal{S}(\text{SDP-PEP})}\mathbf{F}\vec{\mathbf{f}}_{K},$ (9)
@@
Recall that for FGM, there are two sequences, the primary sequence $\{\mathbf{y}_{k}\}$ and the secondary one $\{\mathbf{x}_{k}\}$. The OGM is optimized for minimizing $f(\mathbf{x}_{k})-f_{*}$ while the commonly used measure is $f(\mathbf{y}_{K})-f_{*}$. Due to this fact, [40] propose a new measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ since for smooth function and with OGM update rule, $f(\mathbf{y}_{K+1})=f(\mathbf{x}_{K}-\frac{1}{L}\nabla f(\mathbf{x}_{K}))\leq f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$. The solution of this new measure results in OGM (Algorithm 4). These two algorithms only differ in the last iteration’s $\alpha_{K}$ update. For simplicity, we only focus on this measure and its corresponding relaxed PEP.
Algorithm 3 OGM with last adjustment
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
4:     $\alpha_{k+1}=\begin{cases}\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2},k\leq K-2\\ \frac{1+\sqrt{1+8\alpha_{k}^{2}}}{2},k=K-1\end{cases}$
5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
6:  end for
Algorithm 4 OGM
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
4:     $\alpha_{k+1}=\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
6:  end for
The relaxed optimization problem is
@@
$=$ $\tau R^{2}+\sum_{k=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
@@
$\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\succeq 0.$
Here we omit the process of relaxing the minimax problem and use the solution in [42] directly to explain the property of this feasible solution with the inexact oracle. Define $A_{k}=\sum_{i=0}^{k}\alpha_{i}$ and $A_{0}=1,A_{-1}=0$, and assume $\alpha_{k}^{2}\leq A_{k}$, the feasible solution of the minimax problem is
@@
The above feasible solution works for both OGM and Generalized OGM (discussed in the next section), which means for any stepsize $\{\alpha_{k}\}$ and $\{A_{k}\}$ that satisfy $\alpha_{k}^{2}\leq A_{k}$, (10) and (11) satisfies the constraints of (D).
It is easy to verify that the equality constraint holds with this solution. We still need to find the feasible solution for $\mathbf{u}$ to satisfy the positive semidefinite (PSD) constraint. After substitution by (10) and (11), the left-hand-side matrix of the PSD constraint is
$\mathbf{M}=\begin{pmatrix}[l]\frac{L}{4A_{K}}&[\mathbf{0}]_{1\times K}&[\mathbf{p}]_{1\times(K+1)}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times(K+1)}\\ [\mathbf{p}^{\top}]_{(K+1)\times 1}&[\mathbf{B}^{\top}]_{(K+1)\times K}&[\mathbf{C}]_{(K+1)\times(K+1)}\end{pmatrix},$
@@
Recalling that $\mathbf{R}^{\backslash[1]}$ denotes the matrix resulting from dropping the first column and row of $\mathbf{R}$, this submatrix is exactly the Shur complement of the block $[\tau]$ of the matrix $\mathbf{M}$. Since $\tau>0$, we know that $\mathbf{M}\succeq 0$ if and only if $\mathbf{R}^{\backslash[1]}\succeq 0$. In the following lemma, we will show that $\mathbf{u}=\infty$ is the unique feasible solution when $A_{i}-\alpha_{i}^{2}=0$.
@@
With the given solution in (10) and (11), and $A_{i}-\alpha_{i}^{2}=0$, $\mathbf{u}=\infty$ is the unique solution that makes $\mathbf{R}^{\backslash[1]}\succeq 0$. When $A_{i}-\alpha_{i}^{2}>0$, $\mathbf{u}=\infty$ is still feasible but not unique.
@@
  1. 1.
    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=0$,
  2. 2.
    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$,
  3. 3.
    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=0$,
  4. 4.
    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$.
For the first and second scenarios, it is obvious that $\mathbf{R}^{\backslash[1]}_{i,j}=0$, with diagonal elements being non-negative, the condition holds. For the third scenario, $\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$. Consider the case $\mathbf{R}^{\backslash[1]}_{i,j}>0$, since $c$ can be any negative number, the condition $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=u_{k}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only when $u_{k}=\infty$. For the fourth scenario, the situation is similar. $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=c^{2}u_{l}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only with $u_{l}=\infty$. We get the final result by combining all these scenarios. ∎
@@ 3 Generalized OGM with inexact gradient oracle
The properties of OGM and its generalization have been thoroughly studied. However, the convergence analysis of OGM with inexact gradient oracle has not been performed yet. Furthermore, the existing analysis of FGM with inexact oracles with the absolute error assumption depends on the trajectory of sequences. This section aims at closing these gaps.
It is possible to optimize the algorithm with the same procedure as OGM, but the solution depends on iteration number $K$, initial condition $R$, and inexactness levels $\{b_{i}\}_{i=1}^{K}$. It is difficult to find a recursive structure of the algorithm that captures the numerical solution (see the experiment results in Appendix B). Hence, we focus on Generalized OGM (GOGM) with inexact oracle and try to find its convergence rate with $b_{i}>0$. Furthermore, the convergence bound of Generalized FGM (GFGM) is also analyzed in Section 3.3.
The GOGM is a generalization of OGM (Algorithm 4) as it only requires the stepsize to satisfy $\alpha_{k}^{2}\leq A_{k}$. The inexact GOGM (iGOGM) is provided in Algorithm 5, in which the gradients in steps 2 and 3 are replaced by their inexact alternatives.
Algorithm 5 Inexact Generalized Optimized Gradient Method (iGOGM)
0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for
The update rule in step 4 guarantees $\alpha_{k}^{2}\leq A_{k}$ when $\lambda_{k}\in[0,1]$ as
$\alpha_{k+1}^{2}=(A_{k+1}-A_{k})^{2}=\lambda_{k+1}A_{k+1}\leq A_{k+1}.$
Furthermore, when $\lambda=1$, the $\mathbf{x}_{k+1}$ update (step 5) reduce to
$\mathbf{x}_{k+1}=\left(1-\frac{1}{\alpha_{k+1}}\right)\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},$
the common update step in both FGM and OGM.
Remark 3.1.
The optimality measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ used in the analysis of exact GOGM is also a meaningful one for the inexact problem. For iGOGM, we have
3.1 PEP approach for the convergence analysis of iGOGM
@@
The goal of this part is to find an analytical feasible solution of (D) with iGOGM stepsize. The feasible solution provides an upper bound of the worst-case convergence rate. To do so, first notice that the values of $\{v_{*,i}\}$ are uniquely determined by $\{v_{i,i+1}\}$ through the equality constraint in (D) and, hence, can be removed from the decision variables. Next, consider the decision variable $\tau$. Recall the objective of the problem (D), $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, contains two components: the first one is the convergence rate and the second one is the accumulated error. Given the convergence rate of the exact GOGM, i.e., $\frac{LR^{2}}{4A_{K}}$, we adjust the objective of (D) as
$\frac{L}{4A_{K}}R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2},$
which is the summation of the convergence rate of exact GOGM and the accumulated error $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$. With these replacements, the decision variables we need to consider are $\{v_{i,i+1}\}$ and $\{u_{i}\}$. To simplify the notation, in the following text, we will use $\mathbf{v}$ and $\mathbf{u}$ to represent the sequences $\{v_{i,i+1}\}$ and $\{u_{i}\}$, respectively.
@@
Solving (D) with this new objective, we observe from the numerical solution that $\mathbf{v}$ is almost equal to its analytical counterpart for problems with exact oracles given in (10). In this experiment, we set $K=\{10,20,\cdots,100\}$, $R=\{1,10,1000\}$, and $L=\{1,10,1000\}$ which result in 90 different scenarios in total. For each scenario, we generate 10 instances with random stepsize parameter $\{\lambda_{k}\}$ and inexactness levels $\{b_{i}\}$. In Figure 1, three examples of the relative difference between the numerical solution of inexact GOGM $v^{num.}_{k}$ and the analytical solution of exact GOGM $v^{anly.}_{k}$ are illustrated. The solid line is an average of 90 instances. Furthermore, Figure 2 illustrates the histogram of relative error for all 900 instances which also supports this conjecture.
@@
As we claimed in Lemma 2.4, OGM with (10) has a unique feasible solution with $\mathbf{u}=\infty$. To make $\mathbf{u}$ to be bounded, we have to guarantee that the stepsize condition holds with strict inequality, i.e. $A_{k}-\alpha_{k}^{2}>0$. Recall the matrix $\mathbf{R}^{\backslash[1]}$
@@
where the elements of the first column of $\mathbf{B}$ and $\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}$ are zero. So we can discard this zero row and column and write the $\mathbf{R}^{\backslash[1,K+2]}$ as
@@
with $\mathbf{D}\triangleq[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]^{\backslash[1]}$.
Example of $\mathbf{R}^{\backslash[1,K+2]}$ with $K=3$ is
$\mathbf{R}^{\backslash[1,K+2]}=\left(\begin{array}[]{ccc|cccc}{u}_{0}&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&{u}_{1}&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{2}&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&\frac{A_{1}-\alpha_{1}^{2}}{LA_{3}}&0&0\\ \frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&0&\frac{A_{2}-\alpha_{2}^{2}}{LA_{3}}&0\\ \frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&0&0&\frac{A_{3}-\alpha_{3}^{2}}{LA_{3}}\end{array}\right).$
We need to find $\mathbf{u}$ such that $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$. Here we notice that $\mathbf{D}=\operatorname{diag}(\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})\succ 0$, hence, it is invertible with all elements strictly bigger than 0. We take the Shur complement of the block $\mathbf{D}$ and define the result matrix as $\mathbf{S}$ with
@@
Example with $K=3$ is
$\mathbf{S}=\begin{pmatrix}{u}_{0}-\frac{\left(1+2\alpha_{1}\right)^{2}}{4LA_{3}\left({A}_{1}-{\alpha}_{1}^{2}\right)}-\frac{{\alpha}_{2}^{2}}{LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{1}-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right)^{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}^{2}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{2}-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right)^{2}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\end{pmatrix}.$
Then, $\mathbf{S}\succeq 0$ iff $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$. A nice property of $\mathbf{S}$ is that $\{u_{i}\}$ only shows in the diagonal elements and all the off-diagonal elements are negative constants. It is known that a symmetric diagonally dominated matrix with non-negative diagonal entries is PSD, with the definition of the diagonally dominated matrix as
@@
Based on the special structure of $\mathbf{S}$, we can simplify the above inequality as
$\mathbf{S}_{i,i}\geq-\sum_{i\neq j}\mathbf{S}_{i,j}.$

### v2

@@ 3.2 PEP for optimization with inexact oracle
and the Gram matrix $\mathbf{G}\triangleq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{S}^{2K+2}$ which is a symmetric positive semidefinite (PSD) matrix. Also, we define sparse vectors $\vec{\mathbf{f}}_{k}\in\mathbb{R}^{K+1}$ and $\vec{\mathbf{x}}_{k},\vec{\mathbf{g}}_{k},\vec{\mathbf{e}}_{k}\in\mathbb{R}^{2K+2}$ to select corresponding columns in $\mathbf{F}$ or $\mathbf{X}$ to recover the needed element, which satisfy
@@
Note. We note that the subscript $k$ on these 0-1 vectors does not mean their $k$-th elements are equal to 1. For instance, based on the definitions of $\mathbf{F}$ and $\mathbf{X}$ in (11), $\vec{\mathbf{f}}_{k}$, $\vec{\mathbf{e}}_{k}$, and $\vec{\mathbf{g}}_{k}$ are standard basis vectors with $k+1$-th, $k+2$-th, $K+k+2$-th elements equal to 1, respectively, and $\vec{\mathbf{x}}_{0}$ is a standard basis vector with its first element equal to 1. With a slight abuse of the notation, $\vec{\mathbf{x}}_{k}$ (which is not a basis vector) is the vector used to represent $\mathbf{x}_{k}$ by a linear combination of the columns of $\mathbf{X}$, and it is a function of $\theta_{k,i}$.
@@
The optimal solution of (SDP-PEP) will provide an exact worst-case performance of a given algorithm for convex problems.
It is possible to find the optimal algorithm with parameter $\boldsymbol{\theta}$} by solving the minimax problem
@@
where $\mathbf{G},\mathbf{F}\in\mathcal{S}(\text{SDP-PEP})$ refers to $\mathbf{G}$ and $\mathbf{F}$ satisfy the constraints of (SDP-PEP). Since $\mathbf{A}^{i,j}$ is indeed a function of $\{\theta_{k,i}\}$, this minimax problem has bilinear terms in the constraints of the inner problem, and solving it for the global solution is NP-hard. OGM’s [18] idea is to select a subset of the constraints and upper bound the inner maximization by its dual minimization problem. By wisely selecting the constraints, the hard minimax problem can be relaxed as a bilinear minimization problem which can be further relaxed as a linear SDP. Solving this final problem numerically inspires a theoretical solution and results in the first version of OGM when gradients are exact.
Recall that for FGM, there are two sequences, the primary sequence $\{\mathbf{y}_{k}\}$ and the secondary one $\{\mathbf{x}_{k}\}$. The OGM is optimized for minimizing $f(\mathbf{x}_{k})-f_{*}$ while the commonly used measure is $f(\mathbf{y}_{K})-f_{*}$. Due to this fact, Kim and Fessler [33] propose a new measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ since for smooth function and with OGM update rule, $f(\mathbf{y}_{K+1})=f(\mathbf{x}_{K}-\frac{1}{L}\nabla f(\mathbf{x}_{K}))\leq f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$.
Based on the constraint selection and objective modification discussed above, the relaxed optimization problem is
@@
$=$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
@@
[…]
Kim and Fessler [32] analytically solved (D) nested inside the minimization over the algorithm parameters $\boldsymbol{\theta}$}, in the absence of gradient inexactness, i.e., when $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}=0$ and $\sum_{i=0}^{K}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}=0$, and derived OGM with the following solution,
@@
with $A_{k}=\sum_{i=0}^{k}\alpha_{i}$, $A_{0}=1,A_{-1}=0$. Further, they showed that this solution is also a feasible solution to GOGM [35] with $\alpha_{k}^{2}\leq A_{k}$. We will follow this process, use the feasible solution above, and derive the value of $\mathbf{u}$ for GOGM with the inexact gradients.
@@
We attempted to derive an optimized inexact first-order method by extending the procedure used for OGM [32]. However, this approach revealed a significant practical limitation. The core issue is that the derived optimal step sizes for the inexact method lack a recursive structure. As shown in Appendix B, in the exact gradient case, the step size for a given iteration is independent of the total number of iterations $K$. This allows for a practical, recursive implementation where past iterates remain fixed. In the inexact setting, however, this property does not hold. The optimal step sizes for all previous iterations depend on the pre-specified total number of iterations $K$. This means that if one were to extend a run by increasing $K$, all previously computed steps would no longer be optimal and would need to be recomputed. This phenomenon makes such an algorithm computationally impractical for implementation.
It is easy to verify that the equality constraint in (D) holds with this solution. We still need to find a feasible solution for $\mathbf{u}$ to satisfy the PSD constraint in (14). After substitution by (15) and (16), the left-hand-side matrix of the PSD constraint (denoted as $\mathbf{M}$) in (14) is
$\mathbf{M}=\begin{pmatrix}[l]\tau&[\mathbf{0}]_{1\times K}&[\mathbf{p}]_{1\times(K+1)}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times(K+1)}\\ [\mathbf{p}^{\top}]_{(K+1)\times 1}&[\mathbf{B}^{\top}]_{(K+1)\times K}&[\mathbf{C}]_{(K+1)\times(K+1)}\end{pmatrix}$,
@@
Recalling that $\mathbf{R}^{\backslash[1]}$ denotes the matrix resulting from dropping the first column and row of $\mathbf{R}$, this submatrix is exactly the Shur complement of the block $\tau]$ of the matrix $\mathbf{M}$. Since $\tau>0$ (as $\tau=\frac{L}{4A_{K}}$), we know that $\mathbf{M}\succeq 0$ if and only if $\mathbf{R}^{\backslash[1]}\succeq 0$. In the following lemma, we will show that $\mathbf{u}=\infty$ is the unique feasible solution when $A_{i}=\alpha_{i}^{2},\forall i=1,\cdots,K$.
@@
With the given solution in (15) and (16), if $A_{i}-\alpha_{i}^{2}=0,\forall i=1,\cdots,K$, then $\mathbf{u}=\infty$ is the unique solution that makes $\mathbf{R}^{\backslash[1]}\succeq 0$. When $A_{i}-\alpha_{i}^{2}>0$, $\mathbf{u}=\infty$ is still a feasible solution but not the unique one.
@@
[…]
For the first and second scenarios, it is obvious that $\mathbf{R}^{\backslash[1]}_{i,j}=0$, with diagonal elements being non-negative, the condition holds. For the third scenario, $\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$. Consider the case $\mathbf{R}^{\backslash[1]}_{i,j}>0$, since $c$ can be any negative number, the condition $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=u_{k}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only when $u_{k}=\infty$. For the fourth scenario, the situation is similar. $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=c^{2}u_{l}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only with $u_{l}=\infty$. ∎
@@ 4 Generalized OGM with inexact gradient oracle
The goal of this part is to find an analytical feasible solution of (D) with the iGOGM stepsize. From the weak duality theorem, the feasible solution provides an upper bound on the worst-case convergence rate. To do so, first notice that the values of $\{v_{*,i}\}$ are uniquely determined by $\{v_{i,i+1}\}$ through the equality constraint in (D) and, hence, can be removed from the decision variables. Next, consider the decision variable $\tau$. Recall the objective of the problem (D), $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, contains two components: the first one is the convergence rate and the second one is the accumulated error. Given the convergence rate of the exact GOGM, i.e., $\frac{LR^{2}}{4A_{K}}$, we adjust the objective of (D) as $\frac{L}{4A_{K}}R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, which is the summation of the convergence rate of exact GOGM and the accumulated error $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$. With these replacements, the decision variables we need to consider are $\{v_{i,i+1}\}$ and $\{u_{i}\}$. To simplify the notation, below, we will use $\mathbf{v}$ and $\mathbf{u}$ to represent the sequences $\{v_{i,i+1}\}$ and $\{u_{i}\}$, respectively.
@@
Solving (D) with the new objective, we observe from the numerical solution that $\mathbf{v}$ is almost equal to its analytical counterpart for problems with exact oracles given in (15). In this experiment, we set $K=\{10,20,\cdots,100\}$, $R=\{1,10,1000\}$, and $L=\{1,10,1000\}$ which result in 90 different scenarios in total. For each scenario, we generate 10 instances with random stepsize parameter $\{\lambda_{k}\}$ and inexactness levels $\{b_{i}\}$. In Figure 5, three examples of the relative difference between the numerical solution of iGOGM $v^{num.}_{k}$ and the analytical solution of exact GOGM $v^{anly.}_{k}$ are illustrated. The solid line is an average of 90 instances. Furthermore, Figure 6 illustrates the histogram of relative error for all 900 instances which also supports this conjecture.
@@
As we claimed in Lemma 3.5, OGM ($A_{k}=\alpha_{k}^{2}$) with (15) has a unique feasible solution with $\mathbf{u}=\infty$. To make $\mathbf{u}$ to be bounded, we have to guarantee that the stepsize condition holds with strict inequality, i.e., $A_{k}-\alpha_{k}^{2}>0$. Recall the matrix $\mathbf{R}^{\backslash[1]}$
@@
where the elements of the first column of $\mathbf{B}$ and $\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}$ are zero. So we can discard this zero row and column and write the resulting matrix $\mathbf{R}^{\backslash[1,K+2]}$ as
@@
with $\mathbf{D}\triangleq[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]^{\backslash[1]}$. The example of $\mathbf{R}^{\backslash[1,K+2]}$ with $K=3$ is
[…]
We need to find $\mathbf{u}$ such that $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$. Here we notice that $\mathbf{D}=\operatorname{diag}(\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})$, hence, it is invertible as all diagonal elements are strictly bigger than 0. We take the Shur complement of the block $\mathbf{D}$ and define the result matrix as $\mathbf{S}$ where
@@
The example with $K=3$ is
[…]
Then, we have $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$ iff $\mathbf{S}\succeq 0$. A nice property of $\mathbf{S}$ is that $\{u_{i}\}$ only shows in the diagonal elements, and all the off-diagonal elements are negative. To ensure $\mathbf{S}\succeq 0$, we use the property of the diagonally dominated matrix, which is that a symmetric diagonally dominated matrix with non-negative diagonal entries is PSD, with the definition of the diagonally dominated matrix as
@@
Based on the special structure of $\mathbf{S}$, we can simplify the above inequality as $\mathbf{S}_{i,i}\geq-\sum_{i\neq j}\mathbf{S}_{i,j}$.